OpenStudy (steven513):

The maximum horizontal range of a projectile is given by the formula R=u^2/g, where u is the initial velocity and g is the acceleration due to gravity. Find the velocity with which a ball can be thrown to have a maximum range of 20 meters when the acceleration due to gravity is equal to 9.8 m/s. My answer: 20=u^2/9.8 118=u^2 u=10.8 Am I correct? Thanks For Checking

OpenStudy (sshayer):

\[R=\frac{ u^2 }{ g },u^2=R*g\] \[u^2=20*9.8=196,u=\sqrt{196}=14m/s\]